Calculate v(t) and a of Functions Graph using Matlab xe), vC), a(t)
2ttt cox 2. lOt-4t+2¢? t2(10t4t t3 1 +0 -t dt 1e -t
Calculate v(t) and a of Functions Graph using Matlab xe), vC), a(t)
2ttt cox 2. lOt"-4t+2¢? t2(10t4t t3 1 +0 -t dt 1e -t
Answer #1

MATLAB CODE:

clc
clear all
close all

syms t

x(t) = t^3 - t^4 + 2*t^5; % Distance
v(t) = diff(x(t),t,1);    % Velocity
a(t) = diff(x(t),t,2);    % Acceleration

fprintf('x(t) = ')
disp(x(t))

fprintf('v(t) = ')
disp(v(t))

fprintf('a(t) = ')
disp(a(t))

subplot(3,1,1)
fplot(t,x(t),'k','LineWidth',2)
xlabel('t')
title('x(t)')
grid on

subplot(3,1,2)
fplot(t,v(t),'k','LineWidth',2)
xlabel('t')
title('v(t)')
grid on

subplot(3,1,3)
fplot(t,a(t),'k','LineWidth',2)
xlabel('t')
title('a(t)')
grid on

OUTPUT:

2*t 5 t^4 t^3 x (t) 10*t 4 4*t 3 3*t^2 v (t) 40*t 3 12*t 2 6*t a (t)

x(t) 4000 2000 -2000 -4000 -6000 5 2 1 1 2 4 5 v(t) 6000 5000 4000 3000 2000 1000 5 2 1 2 5 t a(t) 4000 2000 -2000 -4000 5 -2